For how long must an object travel at 29.8km/s to cover 940Gm?
1
Expert's answer
2019-09-13T10:15:32-0400

"t = \frac{940\text{ Gm}}{29.8\text{ km/s}} = \frac{940 \cdot 10^9 \text{ m}}{29.8 \cdot 10^3 \text{ m}} \text{ s} = \frac{9.4 \cdot 10^{11} \text{ m}}{2.98 \cdot 10^4 \text{ m}} \text{ s} \approx 3.154 \cdot 10^7 \text{ s}."

To have better understanding of the magnitude of this number, we may convert remembering that there is 86400 seconds in 1 day (i.e. 24 hours):

"t = \frac{3.15 \cdot 10^7 \text{ s}}{86400\text{ s/days}} \approx 365\text{ days} \approx 1\text{ year}."